OpenStudy (faiqraees):

The sequence a1,a1,a3 is such that \(a_n>5\) and \(\a_{n+1} = \frac{4a_n}{5}+\frac{5}{a_n}\) for every positive integer n. Prove by mathematical induction that \(a_n>5\) for every positive integer n.

OpenStudy (faiqraees):

How am I supposed to prove for a(k+1)?

OpenStudy (math&ing001):

suppose it's true for a(k) then prove it's true for a(k+1)

OpenStudy (karim728):

I think this is a reverse of mathematical induction

OpenStudy (faiqraees):

I am well aware of the method. I can't however get a conclusion for a(k+1)

OpenStudy (faiqraees):

I can get to the form \[a_{k+1}> \frac{25}{a_n} \] From there, I dont know how to prove \(a_{k+1}>5\)

OpenStudy (faiqraees):

My method: \(a_k>5\) \((a_k)^2>25\) \(4(a_k)^2>100\) \(4(a_k)^2+25>125\) \[\frac{4(a_k)^2+25}{5a_n}>\frac{25}{a_n}\] \(a_{k+1}>\frac{25}{a_n}\) Not sure what to do next

OpenStudy (faiqraees):

@ParthKohli

OpenStudy (math&ing001):

I found the proof, but in kind of a dirty way. You analyze it as a function f(x) = (4x^2+25)/5x You'll find that the function is strictly increasing for x>5 and that the minimum value at x=5 is f(x)=5. Meaning for x>5, f(x)>5 So for a_(k) > 5, a_(k+1)>5

OpenStudy (faiqraees):

\(\color{#0cbb34}{\text{Originally Posted by}}\) @math&ing001 I found the proof, but in kind of a dirty way. You analyze it as a function f(x) = (4x^2+25)/5x You'll find that the function is strictly increasing for x>5 and that the minimum value at x=5 is f(x)=5. Meaning for x>5, f(x)>5 So for a_(k) > 5, a_(k+1)>5 \(\color{#0cbb34}{\text{End of Quote}}\) Hats off to your brilliance. Thanks

OpenStudy (math&ing001):

Welcome !

OpenStudy (eliesaab):

Here is the same solution that I was going to post. I think that I posted for another problem before \[ g(x)=\frac{4 x}{5}+\frac{5}{x}\\ g'(x)=\frac{4}{5}-\frac{5}{x^2}=\frac{(2 x-5) (2 x+5)}{5 x^2} \] so g is increasing as x>2,5 and g(2.5)=5 All the a_n are bigger than at least 4 We are done